solve
having problems with this
d(d-3)=40
show work

Answer by tutorcecilia(2152)

d(d-3)=40
First, simply by using the distributive property:
d^2 – 3d =40
.
Since the highest degree is 2 (d^2) try to set it up as a quadratic polynomial. Set the entire equation equal to zero by subtracting 40 from both sides.
d^2 – 3d - 40=0
Factor:
(d+5)(d-8)=0
Set each factor equal to zero and solve for d:
d+5= 0
d = -5
.
d-8=0
d =8
The possible solution set is 3 and/or 8.
Check each solution by plugging it back into the original problem:
d(d-3)=40
-5(-5-3)=40
-5(-8)=40
40=40.
.
d(d-3)=40
8(8-3)=40
8(5)=40
40=40